OpenStudy (anonymous):

Can someone check? Solve expression using properties of logarithms. 1/2log (basex) – log(basex) = log (base2) = 1/4?

OpenStudy (shubhamsrg):

?? please clarify the ques..

OpenStudy (anonymous):

Use properties of logarithms to solve the equations is the answer 1/4

OpenStudy (shubhamsrg):

1/2log(base x) ,,what does this mean ?

OpenStudy (anonymous):

so it would be like 1/2 log x and x would be the base

OpenStudy (shubhamsrg):

you mean log (1/2) ?? with base x ?

OpenStudy (anonymous):

that would be mean that 1/2 is the base

OpenStudy (shubhamsrg):

ohh..i see.. and what does log(basex) mean?

OpenStudy (anonymous):

its the base

OpenStudy (shubhamsrg):

of what ?

OpenStudy (anonymous):

thst specific log

OpenStudy (shubhamsrg):

log x of what base ??

OpenStudy (anonymous):

x. honestly im a bit confused myself.

mathslover (mathslover):

is this the question?

OpenStudy (shubhamsrg):

lol..

mathslover (mathslover):

@vishweshshrimali5

OpenStudy (vishweshshrimali5):

OMG a lot confusing I am sure there is some mistake

OpenStudy (vishweshshrimali5):

@shubhamsrg what do u think.............. ? what should we do with only base of log

OpenStudy (shubhamsrg):

this is abusrd,,the question has no meaning..

OpenStudy (anonymous):

lol

OpenStudy (vishweshshrimali5):

yes......... logs with only bases have no meaning lol

OpenStudy (anonymous):

@mathslover yes that is the question!

OpenStudy (anonymous):

logs with no bases are by default a base of 10

mathslover (mathslover):

haha

mathslover (mathslover):

Bagatrix says that question is right hahahahahah

OpenStudy (vishweshshrimali5):

but what if logs have ONLY bases

OpenStudy (anonymous):

yes.. but is it assumed that these equal 1 if they only have a base?

OpenStudy (anonymous):

the base is x lol

mathslover (mathslover):

I don't know abt that but the question is true :D

mathslover (mathslover):

Please don't blame the asker hhaah

OpenStudy (anonymous):

wait where is it?

OpenStudy (shubhamsrg):

ofcorse the ques is false!! if log(base2) = y ,say then 2^y = ?? you need to have a quantity there!!

OpenStudy (anonymous):

would it be 1/2log(base x)1-log(base x)1= etc...

OpenStudy (vishweshshrimali5):

2^y = ?? multiply by 0 both sides, 0 = 0 which is true hence the statement is true

OpenStudy (vishweshshrimali5):

lollllllllllll

hartnn (hartnn):

read the question ignoring the word base....everything will make sense..... there is printing mistake which made x as base....its not....its log x ,base unspecified

OpenStudy (shubhamsrg):

thank god you said "lolll" ,,otherwise i was taking your comment seriously and was going to pounce on you! :P

OpenStudy (vishweshshrimali5):

really ??

OpenStudy (anonymous):

one does not simply ignore words

OpenStudy (vishweshshrimali5):

yes.......

mathslover (mathslover):

But what abt my document?

mathslover (mathslover):

:D

hartnn (hartnn):

 yes you are correct :)